# ICM-Assistant: Instruction-tuning Multimodal Large Language Models for Rule-based Explainable Image Content Moderation

### v1

@@ Abstract
Controversial contents largely inundate the Internet, infringing various cultural norms and child protection standards. Traditional Image Content Moderation (ICM) models fall short in producing precise moderation decisions for diverse standards, while recent multimodal large language models (MLLMs), when adopted to general rule-based ICM, often produce classification and explanation results that are inconsistent with human moderators. Aiming at flexible, explainable, and accurate ICM, we design a novel rule-based dataset generation pipeline, decomposing concise human-defined rules and leveraging well-designed multi-stage prompts to enrich short explicit image annotations. Our ICM-Instruct dataset includes detailed moderation explanation and moderation Q-A pairs. Built upon it, we create our ICM-Assistant model in the framework of rule-based ICM, making it readily applicable in real practice. Our ICM-Assistant model demonstrates exceptional performance and flexibility. Specifically, it significantly outperforms existing approaches on various sources, improving both the moderation classification (36.8% on average) and moderation explanation quality (26.6% on average) consistently over existing MLLMs. Code/Data is available at https://github.com/zhaoyuzhi/ICM-Assistant.
@@
footnotetext: Internship with Huawei Hong Kong Research Center.footnotetext: Corresponding Author and Project Lead.**footnotetext: Equal Contribution.
@@ Introduction
There are three technical contributions: (1) The flexible dataset generation pipeline based on concise moderation rules, short human annotations, and well-designed multi-stage prompts for data augmentation, largely surpassing existing moderation datasets; (2) The large-scale instruction-tuning ICM-Instruct dataset and an efficient evaluation benchmark for multiple ICM tasks; (3) The first method to accomplish diverse content moderation tasks, including moderation classification, explanation, and question-answering by adopting several instruction-tuning strategies to upgrade several MLLMs with ICM-Assistant. Results demonstrate that our method well moderates a wide range of images, outperforming existing MLLMs by an average of 33.7% in moderation classification and 26.6% in moderation explanation on various moderation tasks, rules, and data sources (e.g., user-generated, AI-generated, etc.).
@@
Figure 2: The overall pipeline from a specific set of moderation rules to a rule-based ICM-Assistant model. The first row illustrates the procedure of rule decomposition, image downloading, and initialization of explicit descriptions. The second row presents the fully automatic data augmentation pipeline for ICM-Instruct dataset and the instruction-tuning process for ICM-Assistant models.
@@
Figure 3: Pipeline for generating moderation explanations (left) and Q-A pairs (right), from one attribute products “lower-leg”.
@@
Observing the human moderators, they achieve proper moderation usually by first describing the images and then locating regions/semantics that violate the moderation rules. To reduce the workload and achieve flexible moderation, we start with a rule decomposition pipeline to split the rule items into short annotations for collecting images and writing explicit descriptions (one sentence or phrase related to the moderation rules) by moderators. Based on our observations, the rules defined for a moderation term (e.g., “sexy”) can be decomposed into multiple orthogonal basis vectors, terms as attributes. The values assigned to each orthogonal attribute are deliberately chosen to be distinctive in semantics (e.g., lower, middle, and higher in position), such that all the rule items can be represented as products of such attributes are distinct, termed attributes products.
To better explain, we use “sexy” as an example moderation term (Search 2024; Facebook 2024) to illustrate (Fig. 2 (top)): (1) images with sexy attraction content (e.g., a person’s actions and clothing) are considered violations. Our method thereby decomposes the moderation rule items for “sexy” on the aspects of body bareness, actions of people, and camera focus position on the image. We then retrieve attribute products to summarize possible violations in the rule items, e.g, “lower leg” and “upper take off.” Based on the attribute products, a dataset with real images is downloaded from the Internet by searching them on multiple search engines, including Google, Baidu, and Bing. A set of explicit descriptions paired with moderation classification results for each attribute product is also prepared by human moderators as well (see Fig. 2 (top right); in our implementation, there are 35 short sentences in total for the moderation term “sexy,” following an example rule). More details can be found in Supp. A.
@@
Figure 4: Illustration of process for building ICM-Test dataset.
@@
The validation set includes 1,623 images, 1,623 classification labels, 379 moderation explanations from human moderators (a random subset of 1,623 images), and 17,701 moderation Q-A, based on rules R1/R2, termed separately as ICM-Val-R1 and ICM-Val-R2. In addition, there are two test sets adopted in our experiments for further evaluation of the moderation ability for S-ICM, termed ICM-Test, which includes user-generated images and AI-generated images as follows: (1) UGC: 10,740 user-uploaded advertisement images downloaded from diverse advertisement platforms, where there are 1,443 not-sexy images, labeled according to S-ICM R1. (See Fig. 4 (a).) We keep the data ratio in our test set to simulate real-world scenarios; (2) AIGC: 13,140 images generated by SDXL-1.0 (Podell et al. 2023). There are 1,000 initial SDXL-styled prompts automatically generated according to S-ICM R1. We generate 1,140 not-sexy images and 12,000 sexy images. (see Fig. 4 (b)) Human moderators have carefully checked all the above data samples and labels, and there are no privacy issues, as all the exposed real faces are blurred. Besides the ICM-Instruct for S-ICM, we collect relatively small sets for moderation terms “horrifying” and “gambling” for experiments. We provided more details in Supp. A.
@@ I M C E Q
Table 2: Results on ICM accuracy. Our ICM-assistant consistently brings signifcant improvement on four different MLLMs for both the ICM-Val sets on rules R1 and R2.
@@
Figure 6: Comparing the moderation quality of ICM-LLaVA-v1.5 with its base model (marked in gray and yellow separately). Green texts indicate consistent replies, whereas red texts indicate inconsistent replies.
@@
Moderation explanation evaluation is based on MEQ and MQA acc. From Table 4, we notice the clear advantages of our ICM-Assitant models. Compared with the pre-trained MLLMs, our model achieves higher scores, with 26.6% higher in MEQ and 17.8% higher in MQA acc. on average. Fig. 6 shows examples of moderation explanation of LLaVA-v1.5 and our ICM-LLaVA-v1.5. For LLaVA-v1.5, Fig. 6 (top) shows its inconsistency in both classification and moderation explanation results, by missing the possibility of unseen sexual acts. In Fig. 6 (bottom), though the image is considered to have a sexual act, the shape, appearance, and position of the orange and cherry are missed in deducting the reasons. Our method can provide consistent answers with those from human moderators and explain well. More moderation explanation and moderation Q-A results are presented in Supp. C.
@@ Conclusion
In this work, we propose a flexible, explainable, and accurate ICM approach, by building a novel rule-based data generation pipeline, constructing the ICM-Instruct dataset, and instruct-tuning several MLLMs as ICM-Assistants for diverse ICM tasks. Experimental results show effective moderation of web images, user-generated ads, and AI-generated images under different rules. Notably, high accuracy in ICM tasks highlights our approach’s potential in real-world applications, and zero-shot capabilities suggest strong generalizability in other moderation tasks.

### v2

@@ Abstract
Controversial contents largely inundate the Internet, infringing various cultural norms and child protection standards. Traditional Image Content Moderation (ICM) models fall short in producing precise moderation decisions for diverse standards, while recent multimodal large language models (MLLMs), when adopted to general rule-based ICM, often produce classification and explanation results that are inconsistent with human moderators. Aiming at flexible, explainable, and accurate ICM, we design a novel rule-based dataset generation pipeline, decomposing concise human-defined rules and leveraging well-designed multi-stage prompts to enrich short explicit image annotations. Our ICM-Instruct dataset includes detailed moderation explanation and moderation Q-A pairs. Built upon it, we create our ICM-Assistant model in the framework of rule-based ICM, making it readily applicable in real practice. Our ICM-Assistant model demonstrates exceptional performance and flexibility. Specifically, it significantly outperforms existing approaches on various sources, improving both the moderation classification (36.8% on average) and moderation explanation quality (26.6% on average) consistently over existing MLLMs.
@@
Codehttps://github.com/zhaoyuzhi/ICM-Assistant
@@ Introduction
Figure 2: The overall pipeline from a specific set of moderation rules to a rule-based ICM-Assistant model. The first row illustrates rule decomposition, image downloading, and initialization of explicit descriptions. The second row presents the fully automatic data augmentation pipeline for ICM-Instruct dataset and the instruction-tuning process for ICM-Assistant models.
@@
There are three technical contributions in this work:
(1) The flexible dataset generation pipeline based on concise moderation rules, short human annotations, and well-designed multi-stage prompts for prompt-guided data augmentation, largely surpassing existing moderation datasets;
(2) A large-scale instruction-tuning ICM-Instruct dataset and an efficient evaluation benchmark for both the proposed benchmark and existing image moderation test set;
(3) Built upon the ICM-Instruct dataset, we produce the first method to accomplish diverse content moderation tasks, including moderation classification, explanation, and question-answering by adopting several instruction-tuning strategies to upgrade several MLLMs with ICM-Assistant.
Results demonstrate that our method well moderates a wide range of images, outperforming existing MLLMs by an average of 33.7% in moderation classification and 26.6% in moderation explanation on various moderation tasks, rules, and data sources (e.g., user-generated, AI-generated, etc.).
@@
Figure 3: Pipeline for generating moderation explanations and Q-A pairs, from one attribute products “lower-leg”.
@@
Observing the human moderators, they usually achieve proper moderation by first describing the images and then locating regions/semantics that violate the moderation rules. To reduce the workload and achieve flexible moderation, we start with a rule decomposition pipeline to split the rule items into short annotations for collecting images and writing explicit descriptions (one sentence or phrase related to the moderation rules) by moderators. Based on our observations, the rules defined for a moderation term (e.g., “sexy”) can be decomposed into multiple orthogonal basis vectors, terms as attributes. The values assigned to each orthogonal attribute are deliberately chosen to be distinctive in semantics (e.g., lower, middle, and higher in position), such that all the rule items can be represented as products of such attributes are distinct, termed attributes products.
To better explain, we use “sexy” as an example moderation term (Search 2024; Facebook 2024) to illustrate (Fig. 2 (top)): (1) images with sexy attraction content (e.g., a person’s actions and clothing) are considered violations. Our method thereby decomposes the moderation rule items for “sexy” on the aspects of body bareness, actions of people, and camera focus position on the image. We then retrieve attribute products to summarize possible violations in the rule items, e.g, “lower leg” and “upper take off.” Based on the attribute products, a dataset with real images is downloaded from the Internet by searching them on multiple search engines, including Google, Baidu, and Bing. A set of explicit descriptions paired with moderation classification results for each attribute product is also prepared by human moderators as well (see Fig. 2 (top right) and Supp. A; in our implementation, there are 35 short sentences in total for the moderation term “sexy,” following an example rule).
@@
The validation set includes 1,623 images, 1,623 classification labels, 379 moderation explanations from human moderators (a random subset of 1,623 images), and 17,701 moderation Q-A, based on rules R1/R2, termed separately as ICM-Val-R1 and ICM-Val-R2. In addition, there are two test sets adopted in our experiments for further evaluation of the moderation ability for S-ICM, termed ICM-Test, which includes user-generated images and AI-generated images as follows: (1) UGC: 10,740 user-uploaded advertisement images downloaded from diverse advertisement platforms, where there are 1,443 not-sexy images, labeled according to S-ICM R1. (See Fig. 4 (a).) We keep the data ratio in our test set to simulate real-world scenarios; (2) AIGC: 13,140 images generated by SDXL-1.0 (Podell et al. 2023). There are 1,000 initial SDXL-styled prompts automatically generated according to S-ICM R1. We generate 1,140 not-sexy images and 12,000 sexy images. (see Fig. 4 (b)) Human moderators have carefully checked all the above data samples and labels, and there are no privacy issues, as all the exposed real faces are blurred. Besides the ICM-Instruct for S-ICM, we collect relatively small sets for moderation terms “horrifying” and “gambling” for experiments (see Supp. A).
@@
Figure 4: Illustration of the process for building ICM-Test dataset, including ICM-Test-UGC (left) and ICM-Test-AIGC (right).
@@ I M C E Q
Table 2: Results on ICM accuracy. Our ICM-Assistant consistently brings significant improvement on four different MLLMs for both the ICM-Val sets on rules R1 and R2.
@@
Figure 6: Comparing the moderation quality of ICM-LLaVA-v1.5 with its base model (marked in gray and lime green separately). Green texts indicate consistent replies, whereas red texts indicate inconsistent replies.
@@
Moderation explanation evaluation is based on MEQ and MQA acc. From Table 4, we notice the clear advantages of our ICM-Assitant models. Compared with the pre-trained MLLMs, our model achieves higher scores, with 26.6% higher in MEQ and 17.8% higher in MQA acc. on average. Fig. 6 shows examples of moderation explanation of LLaVA-v1.5 and our ICM-LLaVA-v1.5. For LLaVA-v1.5, Fig. 6 (left) shows its inconsistency in both classification and moderation explanation results, by missing the possibility of unseen sexual acts. In Fig. 6 (right), though the image is considered to have a sexual act, the shape, appearance, and position of the orange and cherry are missed by LLaVA-v1.5 in deducting the reasons. Our method can provide consistent answers with those from human moderators and explain well. More moderation explanation and Q-A results are presented in Supp. C.
@@ Conclusion
In this work, we propose a flexible, explainable, and accurate ICM approach by introducing a novel moderation rule-based data generation pipeline, constructing the ICM-Instruct dataset, and creating several instruction-tuned MLLMs as ICM-Assistants to accomplish diverse explainable content moderation tasks. Experimental results show that our approach moderates web-collected images, user-generated advertisements, and AI-generated images with multiple rules. Meanwhile, the zero-shot ability in other moderation terms indicates the possibility of extending the moderation tasks as a general ability for MLLMs. The balance of the dataset between different moderation terms and multi-task learning in instruction tuning will be our future works.
Acknowledgments
The work is supported by the National Natural Science Foundation of China (No. 62306220) and the RGC/GRF grant 16206524.